• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Bo ( RC )

    Log entry time 08:33:50 on October 01, 2009

    Entry number 292545

    keyword=Experiment update (8:30am 10/01/09)

    Hall-C has finished their calibration and MCC will restore the beam
    energy to the normal. MCC also want to do some 31 MHz test, but this will
    depend on how the installation is going. If not finished by 10:00am, this
    test will be postponed to next beam study. MCC will start to restore the
    beam to Hall-A. Beam will be sent to the Hall around noon if everything
    goes well.
    Jack will ramp up all the Qs magnets and Dipole magnet in R-arm soon.
    Other than that, detectors are in good shape. We took several pedestal
    runs with Happex DAQ. Also run spot++ in both arms and look fine.